OpenStudy (anonymous):

find the derivative of the function: y=e^(5(x)cos(x))

OpenStudy (anonymous):

\[e ^{5x \cos(x)}\]

OpenStudy (anonymous):

please help...i got this wrong...

OpenStudy (anonymous):

dy/dx= e^5xcosx * 5* (cosx - xsinx)

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

thts the answer dy 5(cosx -xsinx) -- = e^5xcosx---------- dx 1

OpenStudy (anonymous):

look malik diffrentiation e^b = e^b ( db/dx) u just do tht and u will get the answer

OpenStudy (anonymous):

ok thanks